Selected sales and operating data for three divisions of different structural engineering firms are given as follows:

    Division A   Division B   Division C
Sales   $   15,250,000      $   35,250,000      $   25,250,000     
Average operating assets   $   3,050,000      $   7,050,000      $   5,050,000     
Net operating income   $   655,750      $   528,750      $   732,250     
Minimum required rate of return      9.00   %      9.50   %      14.50   %

Required:
1. Compute the return on investment (ROI) for each division using the formula stated in terms of margin and turnover.
2. Compute the residual income (loss) for each division.
3. Assume that each division is presented with an investment opportunity that would yield a 10% rate of return.
a. If performance is being measured by ROI, which division or divisions will probably accept or reject the opportunity?
b. If performance is being measured by residual income, which division or divisions will probably accept or reject the opportunity
Answer #1

a) Calculate rate of return

Division A Division B Division C
Margin 655750/15250000 = 4.3% 528750/35250000 = 1.5% 732250/25250000 = 2.9%
Turnover 15250000/3050000 = 5 35250000/7050000 = 5 25250000/5050000 = 5
ROI 4.3*5 = 21.5% 1.5*5 = 7.5% 2.9*5 = 14.5%

b) Residual income

Division A Division B Division C
Minimum operating income 3050000*9% = 274500 7050000*9.5% = 669750 5050000*14.5% = 732250
Actual operating income 655750 528750 732250
Residual income 381250 -141000 0

3a)

Division A Accept
Division B Reject
Division C Accept

3b)

Division A Accept
Division B Reject
Division C Accept/reject